OpenStudy (anonymous):

John is 4 inches taller than Stephen, who is p inches tall. Which expression shows John's height in inches? 1.p-4 2.p/4 3. p+4 ?

Directrix (directrix):

p+4 J = S + 4 J = p + 4

OpenStudy (anonymous):

thaks :)

Directrix (directrix):

If the assistance of others has been helpful, please consider pressing the "Good Answer" button as a token of your appreciation. :)